# GVE-Louvain: Fast Louvain Algorithm for Community Detection in Shared Memory Setting

### v4

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for identifying such divisions is critical in a number of applications, where the size of datasets have reached significant scales. This technical report presents one of the most efficient multicore implementations of the Louvain algorithm, a high quality community detection method. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our Louvain, which we term as GVE-Louvain, outperforms Vite, Grappolo, and NetworKit Louvain by 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × respectively - achieving a processing rate of $560M$ edges/s on a $3.8B$ edge graph. In addition, GVE-Louvain improves performance at an average rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1. Introduction
In recent years, the collection of data and the relationships among them, represented as graphs, have reached unmatched levels. This has necessitated the design of efficient parallel algorithms for community detection on large networks. Existing studies on Louvain propose several optimizations (Rotta and Noack, 2011; Waltman and Eck, 2013; Gach and Hao, 2014; Traag, 2015; Lu et al., 2015; Ryu and Kim, 2016; Ozaki et al., 2016; Naim et al., 2017; Halappanavar et al., 2017; Ghosh et al., 2018b; Traag et al., 2019; Zhang et al., 2021; Shi et al., 2021; You et al., 2022; Aldabobi et al., 2022) and parallelization techniques (Cheong et al., 2013; Wickramaarachchi et al., 2014; Lu et al., 2015; Zeng and Yu, 2015; Que et al., 2015; Naim et al., 2017; Fazlali et al., 2017; Halappanavar et al., 2017; Zeitz, 2017; Ghosh et al., 2018b; Bhowmik and Vadhiyar, 2019; Gheibi et al., 2020; Shi et al., 2021; Bhowmick et al., 2022). Further, significant research effort has been focused on developing efficient parallel implementations of Louvain algorithm for multicore CPUs (Staudt and Meyerhenke, 2015; Staudt et al., 2016; Fazlali et al., 2017; Halappanavar et al., 2017; Qie et al., 2022), GPUs (Naim et al., 2017), CPU-GPU hybrids (Bhowmik and Vadhiyar, 2019; Mohammadi et al., 2020), multi-GPUs (Cheong et al., 2013; Hricik et al., 2020; Chou and Ghosh, 2022; Gawande et al., 2022), and multi-node systems — CPU only (Ghosh et al., 2018b, a; Sattar and Arifuzzaman, 2022) / CPU-GPU hybrids (Bhowmick et al., 2022).
@@
In this report, we introduce our parallel multicore implementation of the Louvain algorithm [^1] https://github.com/puzzlef/louvain-communities-openmp. Our implementation employs asynchronous computation, utilizes parallel prefix sum and preallocated Compressed Sparse Row (CSR) data structures for identifying community vertices and storing the super-vertex graph during the aggregation phase, uses fast collision-free per-thread hash tables for the local-moving and aggregation phases, and incorporates an aggregation tolerance to avoid unnecessary aggregation phases. Additionally, we leverage established techniques such as OpenMP’s dynamic loop schedule, limiting the number of iterations per pass, threshold-scaling optimization, and vertex pruning. For each optimization, we determine optimal parameter settings, if there are any. To the best of our knowledge, our implementation represents the most efficient parallel Louvain algorithm implementation on multicore CPUs. We compare our implementation against other state-of-the-art implementations, including multi-core, multi-node, hybrid GPU, multi-GPU, and multi-node multi-GPU implementations, in Table 1. Both direct and indirect comparisons are provided, with details given in Sections 5.2 and A.1 respectively.
@@ 2. Related work
To parallelize the Louvain algorithm, a number of strategies have been attempted. These include using heuristics to break the sequential barrier (Lu et al., 2015), ordering vertices via graph coloring (Halappanavar et al., 2017), performing iterations asynchronously (Que et al., 2015; Shi et al., 2021), using adaptive parallel thread assignment (Fazlali et al., 2017; Naim et al., 2017; Sattar and Arifuzzaman, 2019; Mohammadi et al., 2020), parallelizing the costly first iteration (Wickramaarachchi et al., 2014), using vector based hashtables (Halappanavar et al., 2017), and using sort-reduce instead of hashing (Cheong et al., 2013). Platforms used range from an AMD multicore system (Fazlali et al., 2017), and Intel’s Knight’s Landing, Haswell (Gheibi et al., 2020), SkylakeX, and Cascade Lake (Hossain and Saule, 2021). Other approaches include the use of MapReduce in a BigData batch processing framework (Zeitz, 2017). It should however be noted though that community detection methods such as the Louvain that rely on modularity maximization are known to suffer from resolution limit problem. This prevents identification of communities of certain sizes (Ghosh et al., 2019).
@@
Bhowmick et al. (Bhowmik and Vadhiyar, 2019) introduce HyDetect, a community detection algorithm designed for hybrid CPU-GPU systems, employing a divide-and-conquer strategy. The algorithm partitions the graph among the CPU and GPU components of a node, facilitating independent community detection utilizing the Louvain algorithm in both segments. Bhowmick et al. (Bhowmick et al., 2022) later introduce a multi-node multi-GPU Louvain community detection algorithm, which involves graph partitioning across multiple nodes, refinement through identification and migration of doubtful vertices between processors, and utilization of a hierarchical merging algorithm ensuring that the merged components can be accommodated within a processor at any given point. Chou and Ghosh (Chou and Ghosh, 2022) present Nido, a batched clustering method for GPUs that leverages a partition-based design, and can process graphs larger than the combined GPU memory of a node. Gawande et al. (Gawande et al., 2022) propose cuVite, an ongoing endeavor focusing on distributed Louvain for heterogeneous systems — building upon their previous work involving parallelizing Louvain for community detection on CPU-only distributed systems. Finally, NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes a parallel implementation of the Louvain algorithm.
@@ 4.1. Optimizations for Louvain algorithm
Our optimizations include using OpenMP’s dynamic loop schedule, limiting the number of iterations per pass to $20$, using a tolerance drop rate of $10$, setting an initial tolerance of $0.01$, using an aggregation tolerance of $0.8$, employing vertex pruning, making use of parallel prefix sum and preallocated Compressed Sparse Row (CSR) data structures for finding community vertices and for storing the super-vertex graph during the aggregation phase, and using fast collision-free per-thread hashtables which are well separated in their memory addresses (Far-KV) for the local-moving and aggregation phases of the algorithm. Details on each of the optimizations is given below.
@@ 4.2. Our optimized Louvain implementation
We now explain the implementation of GVE-Louvain in Algorithms 1, 2, and 3. A flow diagram illustrating the first pass of GVE-Louvain is shown in Figure 4.
@@ 4.2.1. Main step of GVE-Louvain
The main step of GVE-Louvain (louvian() function) is outlined in Algorithm 1. It encompasses initialization, the local-moving phase, and the aggregation phase. Here, the louvain() function takes the input graph $G$, and returns the community membership $C$ for each vertex. In line 15, we first initialize the community membership $C$ of each vertex in $G$, and perform passes of the Louvain algorithm, limited to $MAX\_PASSES$ (lines 16-26). In each pass, we initialize the total edge weight of each vertex $K^{\prime}$, the total edge weight of each community $\Sigma^{\prime}$, and the community membership $C^{\prime}$ of each vertex in the current graph $G^{\prime}$ (line 17); and mark all vertices as unprocessed (line 18).
Next, in line 19, we perform the local-moving phase by calling louvainMove() (Algorithm 2), which optimizes community assignments. If the local-moving phase converged in a single iteration, global convergence is implied and we terminate the passes (line 20). Further, if the drop is community count $|\Gamma|$ is too small, we have reached the point of diminishing returns, and thus stop at the current pass (line 22).
In case convergence has not been achieved, we renumber communities (line 23), update top-level community memberships $C$ with dendrogram lookup (line 24), perform the aggregation phase by calling louvainAggregate() (Algorithm 3), and scale the convergence threshold for subsequent passes, i.e., perform threshold scaling (line 26). The next pass continues in line 16. At the end of all passes, we perform a final update of the top-level community memberships $C$ with dendrogram lookup (line 27), and return the top-level community membership $C$ of each vertex in $G$.
@@
7:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
8:$H_{t}$: Collision-free per-thread hashtable
9:$l_{i}$: Number of iterations performed (per pass)
10:$l_{p}$: Number of passes performed
11:$\tau$: Per iteration tolerance
12:$\tau_{agg}$: Aggregation tolerance
13:
14:function louvain($G$)
15:     Vertex membership: $C\leftarrow[0..|V|)$ ; $G^{\prime}\leftarrow G$
@@
17:         $\Sigma^{\prime}\leftarrow K^{\prime}\leftarrow vertexWeights(G^{\prime})$ ; $C^{\prime}\leftarrow[0..|V^{\prime}|)$
18:         Mark all vertices in $G^{\prime}$ as unprocessed
19:         $l_{i}\leftarrow louvainMove(G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime})$triangleright Alg. 2
20:         if $l_{i}\leq 1$ then break $\triangleright$ Globally converged?
21:         $|\Gamma|,|\Gamma_{old}|\leftarrow$ Number of communities in $C$, $C^{\prime}$
22:         if $|\Gamma|/|\Gamma_{old}|>\tau_{agg}$ then break $\triangleright$ Low shrink?
23:         $C^{\prime}\leftarrow$ Renumber communities in $C^{\prime}$
24:         $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
25:         $G^{\prime}\leftarrow louvainAggregate(G^{\prime},C^{\prime})$triangleright Alg. 3
@@
27:     $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
28:     return $C$
@@
5:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
6:$H_{t}$: Collision-free per-thread hashtable
7:$l_{i}$: Number of iterations performed
8:$\tau$: Per iteration tolerance
9:
10:function louvainMove($G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime}$)
@@
12:         Total delta-modularity per iteration: $\Delta Q\leftarrow 0$
13:         for all unprocessed $i\in V^{\prime}$ in parallel do
14:              Mark $i$ as processed (prune)
15:              $H_{t}\leftarrow scanCommunities(\{\},G^{\prime},C^{\prime},i,false)$
@@
17:              $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$
18:              $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
19:              if $c^{*}=C^{\prime}[i]$ then continue
20:              $\Sigma^{\prime}[C^{\prime}[i]]-=K^{\prime}[i]$ ; $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomic
21:              $C^{\prime}[i]\leftarrow c^{*}$ ; $\Delta Q\leftarrow\Delta Q+\delta Q^{*}$
22:              Mark neighbors of $i$ as unprocessed
23:         if $\Delta Q\leq\tau$ then break $\triangleright$ Locally converged?
24:     return $l_{i}$
25:
26:function scanCommunities($H_{t},G^{\prime},C^{\prime},i,self$)
27:     for all $(j,w)\in G^{\prime}.edges(i)$ do
28:         if $self$ or $i\neq j$ then $H_{t}\leftarrow H_{t}[C^{\prime}[j]]+w$
@@ 4.2.2. Local-moving phase of GVE-Louvain
Lines 11-23 represent the main loop of the local-moving phase. In line 12, we first initialize the total delta-modularity per iteration $\Delta Q$. Next, in lines 22-22, we iterate over unprocessed vertices in parallel. For each vertex $i$, we mark $i$ as processed - vertex pruning (line 14), scan communities connected to $i$ excluding self (line 15), determine the best community $c*$ to move $i$ to (line 17), calculate the delta-modularity of moving $i$ to $c*$ (line 18), and update the community membership (lines 20-21) of $i$, and mark its neighbors as unprocessed (line 22) if a better community was found. In line 23, we check if the local-moving phase has converged (locally). If so, we break out of the loop (or if $MAX\_ITERATIONS$ is reached). At the end, in line 24, we return the number of iterations performed $l_{i}$.

### v5

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for identifying such divisions is critical in a number of applications, where the size of datasets have reached significant scales. This technical report presents one of the most efficient multicore implementations of the Louvain algorithm, a high quality community detection method. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our Louvain, which we term as GVE-Louvain, outperforms Vite, Grappolo, NetworKit Louvain, and cuGraph Louvain (running on NVIDIA A100 GPU) by 50×50$\times$50 ×, $\times2$, 20×20$\times$20 ×, and 5.8×5.8$\times$5.8 × faster respectively - achieving a processing rate of $560M$ edges/s on a $3.8B$ edge graph. In addition, GVE-Louvain improves performance at an average rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1. Introduction
In recent years, the collection of data and the relationships among them, represented as graphs, have reached unmatched levels. This has necessitated the design of efficient parallel algorithms for community detection on large networks. Existing studies on Louvain propose several optimizations (Rotta and Noack, 2011; Waltman and Eck, 2013; Gach and Hao, 2014; Traag, 2015; Lu et al., 2015; Ryu and Kim, 2016; Ozaki et al., 2016; Naim et al., 2017; Halappanavar et al., 2017; Ghosh et al., 2018b; Traag et al., 2019; Zhang et al., 2021; Shi et al., 2021; You et al., 2022; Aldabobi et al., 2022) and parallelization techniques (Cheong et al., 2013; Wickramaarachchi et al., 2014; Lu et al., 2015; Zeng and Yu, 2015; Que et al., 2015; Naim et al., 2017; Fazlali et al., 2017; Halappanavar et al., 2017; Zeitz, 2017; Ghosh et al., 2018b; Bhowmik and Vadhiyar, 2019; Gheibi et al., 2020; Shi et al., 2021; Bhowmick et al., 2022). Further, significant research effort has been focused on developing efficient parallel implementations of Louvain algorithm for multicore CPUs (Staudt and Meyerhenke, 2015; Staudt et al., 2016; Fazlali et al., 2017; Halappanavar et al., 2017; Qie et al., 2022), GPUs (Naim et al., 2017), CPU-GPU hybrids (Bhowmik and Vadhiyar, 2019; Mohammadi et al., 2020), multi-GPUs (Cheong et al., 2013; Kang et al., 2023; Chou and Ghosh, 2022; Gawande et al., 2022), and multi-node systems — CPU only (Ghosh et al., 2018b, a; Sattar and Arifuzzaman, 2022) / CPU-GPU hybrids (Bhowmick et al., 2022).
@@
In this report, we introduce our parallel multicore implementation of the Louvain algorithm [^1] https://github.com/puzzlef/louvain-communities-openmp. Our implementation employs asynchronous computation, utilizes parallel prefix sum and preallocated Compressed Sparse Row (CSR) data structures for identifying community vertices and storing the super-vertex graph during the aggregation phase, uses fast collision-free per-thread hash tables for the local-moving and aggregation phases, and incorporates an aggregation tolerance to avoid unnecessary aggregation phases. Additionally, we leverage established techniques such as threshold-scaling optimization, vertex pruning, limiting the number of iterations per pass, and OpenMP’s dynamic loop schedule. For each optimization, we determine optimal parameter settings, if there are any. To the best of our knowledge, our implementation represents the most efficient parallel Louvain algorithm implementation on multicore CPUs. We compare our implementation against other state-of-the-art implementations, including multi-core, multi-node, hybrid GPU, multi-GPU, and multi-node multi-GPU implementations, in Table 1. Both direct and indirect comparisons are provided, with details given in Sections 5.2 and A.1 respectively.
@@ 2. Related work
To parallelize the Louvain algorithm, a number of strategies have been attempted. These include using heuristics to break the sequential barrier (Lu et al., 2015), ordering vertices via graph coloring (Halappanavar et al., 2017), performing iterations asynchronously (Que et al., 2015; Shi et al., 2021), using adaptive parallel thread assignment (Fazlali et al., 2017; Naim et al., 2017; Sattar and Arifuzzaman, 2019; Mohammadi et al., 2020), parallelizing the costly first iteration (Wickramaarachchi et al., 2014), using vector based hashtables (Halappanavar et al., 2017), and using sort-reduce instead of hashing (Cheong et al., 2013). It should however be noted though that community detection methods that rely on modularity maximization are known to suffer from resolution limit problem (Ghosh et al., 2019).
@@
Bhowmick et al. (Bhowmik and Vadhiyar, 2019) introduce HyDetect, a community detection algorithm designed for hybrid CPU-GPU systems, employing a divide-and-conquer strategy. The algorithm partitions the graph among the CPU and GPU components of a node, facilitating independent community detection utilizing the Louvain algorithm in both segments. Bhowmick et al. (Bhowmick et al., 2022) later introduce a multi-node multi-GPU Louvain community detection algorithm, which involves graph partitioning across multiple nodes, refinement through identification and migration of doubtful vertices between processors, and utilization of a hierarchical merging algorithm ensuring that the merged components can be accommodated within a processor at any given point. Chou and Ghosh (Chou and Ghosh, 2022) present Nido, a batched clustering method for GPUs that leverages a partition-based design, and can process graphs larger than the combined GPU memory of a node. Gawande et al. (Gawande et al., 2022) propose cuVite, an ongoing endeavor focusing on distributed Louvain for heterogeneous systems — building upon their previous work involving parallelizing Louvain for community detection on CPU-only distributed systems. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes a parallel implementation of the Louvain algorithm. Finally, cuGraph (Kang et al., 2023) is a GPU-accelerated graph analytics library that is part of the RAPIDS suite of data science and machine learning tools. It harnesses the power of NVIDIA GPUs to significantly speed up graph analytics compared to traditional CPU-based methods. cuGraph’s core is written in C++ with CUDA and is primarily accessed through a Python interface.
@@ 4.1. Optimizations for Louvain algorithm
Our optimizations include using OpenMP’s dynamic loop schedule, limiting the number of iterations per pass to $20$, using a tolerance drop rate of $10$, setting an initial tolerance of $0.01$, using an aggregation tolerance of $0.8$, employing vertex pruning, making use of parallel prefix sum and preallocated Compressed Sparse Row (CSR) data structures for finding community vertices and for storing the super-vertex graph during the aggregation phase, and using fast collision-free per-thread hashtables which are well separated in their memory addresses (Far-KV) for the local-moving and aggregation phases of the algorithm. Details are given below.
@@ 4.2. Our optimized Louvain implementation
We now explain the implementation of GVE-Louvain in Algorithms 1, 2, and 3. We aim to incorporate GVE-Louvain into our upcoming command-line graph processing tool named ”GVE”, derived from ”Graph(Vertices, Edges)”. GVE-Louvain exhibits a time complexity of $O(KM)$, where $K$ signifies the total iterations conducted. Its space complexity is $O(TN+M)$, where $T$ denotes the number of threads utilized, and $TN$ represents the collision-free hash tables $H_{t}$ employed per thread. A flow diagram illustrating the first pass of GVE-Louvain is shown in Figure 4.
@@ 4.2.1. Main step of GVE-Louvain
The main step of GVE-Louvain (louvian() function) is outlined in Algorithm 1. It encompasses initialization, the local-moving phase, and the aggregation phase. Here, the louvain() function takes the input graph $G$, and returns the community membership $C$ for each vertex. In line 14, we first initialize the community membership $C$ of each vertex in $G$, and perform passes of the Louvain algorithm, limited to $MAX\_PASSES$ (lines 15-25). In each pass, we initialize the total edge weight of each vertex $K^{\prime}$, the total edge weight of each community $\Sigma^{\prime}$, and the community membership $C^{\prime}$ of each vertex in the current graph $G^{\prime}$ (line 16); and mark all vertices as unprocessed (line 17).
@@
7:$H_{t}$: Collision-free per-thread hashtable
8:$l_{i}$: Number of iterations performed (per pass)
9:$l_{p}$: Number of passes performed
10:$\tau$: Per iteration tolerance
11:$\tau_{agg}$: Aggregation tolerance
12:
13:function louvain($G$)
14:     Vertex membership: $C\leftarrow[0..|V|)$ ; $G^{\prime}\leftarrow G$
@@
16:         $\Sigma^{\prime}\leftarrow K^{\prime}\leftarrow vertexWeights(G^{\prime})$ ; $C^{\prime}\leftarrow[0..|V^{\prime}|)$
17:         Mark all vertices in $G^{\prime}$ as unprocessed
18:         $l_{i}\leftarrow louvainMove(G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime})$triangleright Alg. 2
19:         if $l_{i}\leq 1$ then break $\triangleright$ Globally converged?
20:         $|\Gamma|,|\Gamma_{old}|\leftarrow$ Number of communities in $C$, $C^{\prime}$
21:         if $|\Gamma|/|\Gamma_{old}|>\tau_{agg}$ then break $\triangleright$ Low shrink?
22:         $C^{\prime}\leftarrow$ Renumber communities in $C^{\prime}$
23:         $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
24:         $G^{\prime}\leftarrow louvainAggregate(G^{\prime},C^{\prime})$triangleright Alg. 3
@@
26:     $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
27:     return $C$
@@
5:$H_{t}$: Collision-free per-thread hashtable
6:$l_{i}$: Number of iterations performed
7:$\tau$: Per iteration tolerance
8:
9:function louvainMove($G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime}$)
@@
11:         Total delta-modularity per iteration: $\Delta Q\leftarrow 0$
12:         for all unprocessed $i\in V^{\prime}$ in parallel do
13:              Mark $i$ as processed (prune)
14:              $H_{t}\leftarrow scanCommunities(\{\},G^{\prime},C^{\prime},i,false)$
@@
16:              $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$
17:              $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
18:              if $c^{*}=C^{\prime}[i]$ then continue
19:              $\Sigma^{\prime}[C^{\prime}[i]]-=K^{\prime}[i]$ ; $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomic
20:              $C^{\prime}[i]\leftarrow c^{*}$ ; $\Delta Q\leftarrow\Delta Q+\delta Q^{*}$
21:              Mark neighbors of $i$ as unprocessed
22:         if $\Delta Q\leq\tau$ then break $\triangleright$ Locally converged?
23:     return $l_{i}$
24:
25:function scanCommunities($H_{t},G^{\prime},C^{\prime},i,self$)
26:     for all $(j,w)\in G^{\prime}.edges(i)$ do
27:         if not $self$ and $i=j$ then continue
28:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
@@
Next, in line 18, we perform the local-moving phase by calling louvainMove() (Algorithm 2), which optimizes community assignments. If the local-moving phase converged in a single iteration, global convergence is implied and we terminate the passes (line 19). Further, if the drop is community count $|\Gamma|$ is too small, we have reached the point of diminishing returns, and thus stop at the current pass (line 21).
In case convergence has not been achieved, we renumber communities (line 22), update top-level community memberships $C$ with dendrogram lookup (line 23), perform the aggregation phase by calling louvainAggregate() (Algorithm 3), and scale the convergence threshold for subsequent passes, i.e., perform threshold scaling (line 25). The next pass continues in line 15. At the end of all passes, we perform a final update of the top-level community memberships $C$ with dendrogram lookup (line 26), and return the top-level community membership $C$ of each vertex in $G$.
@@ 4.2.2. Local-moving phase of GVE-Louvain
Lines 10-22 represent the main loop of the local-moving phase. In line 11, we first initialize the total delta-modularity per iteration $\Delta Q$. Next, in lines 21-21, we iterate over unprocessed vertices in parallel. For each vertex $i$, we mark $i$ as processed - vertex pruning (line 13), scan communities connected to $i$ excluding self (line 14), determine the best community $c*$ to move $i$ to (line 16), calculate the delta-modularity of moving $i$ to $c*$ (line 17), and update the community membership (lines 19-20) of $i$, and mark its neighbors as unprocessed (line 21) if a better community was found. In line 22, we check if the local-moving phase has converged (locally). If so, we break out of the loop (or if $MAX\_ITERATIONS$ is reached). At the end, in line 23, we return the number of iterations performed $l_{i}$.